The sum of the width and the length of rectangular garden is 103 inches. If the width is 5 inches less than the thrice its length, find the width and the length.
Answer by fcabanski(1391)

The first sentence is l+w=103. The second sentence is w=3l-5. Substitute w from the second equation into the first.


l+3l-5=103 ---> 4l = 108 ---> l = 108/4 = 27.


27+w = 103 ---> w = 103-27 = 76


Check: 76 = 3*27-5 = 81-5 = 76 YES


Contact me for private tutoring: fcabanski@hotmail.com